OpenStudy (anonymous):

Find the quotient: (2x*3 − 3x*2 + x − 6) ÷ (x − 4)

OpenStudy (zehanz):

You could do this by long division or synthetic division - have you heard of (one of) these?

OpenStudy (anonymous):

no i havent

OpenStudy (zehanz):

Synthetic division is the simplest to do. It goes as follows: Write 4 and the coefficients of the 3rd degree polynomial: 4 | 2 -3 1 -6 | |------------- 2 And put the first coefficient (2) straight under the horizontal line. Then add Now multiply 4 by this number 2 and put the result next under the hor line: 4 | 2 -3 1 -6 | 8 |------------- 2 5 Repeat: multiply 4 and 5 and put the result 20 in the next position and add: 4 | 2 -3 1 -6 | 8 20 |------------- 2 5 21 Do this one more time: 4 | 2 -3 1 -6 | 8 20 84 |------------- 2 5 21 78 What you have got now, are the coefficients of a 2nd degree polynomial (2, 5 and 21) and the remainder. In short: \[\frac{ 2x^3 − 3x^2 + x − 6 }{ x-4 }=2x^2+5x+21+\frac{ 78 }{ x-4 }\] At first, it looks complicated, but it is really simple. If you have any questions, just ask!